# A Unified View of Optimal Kernel Hypothesis Testing

### v1

@@ A.2 Proof sketch of differentially private kernel separation
We detail the proof structure of the differentially private kernel separation results: Equation 18 proved in Kim and Schrab (2023, Theorem 7), and Equation 33 proved in Kim and Schrab (2023, Theorem 12). We focus on the V-statistic case $\sqrt{V_{0}}$ with $V_{0}\coloneqq{V_{k}}$ for some kernel $k$, with bootstrapped statistics $\sqrt{V_{1}},\dots,\sqrt{V_{B}}$.
The proof structure is exactly the same as outlined in Section A.1 but taking into account the Laplace privatisation noise $\zeta_{0},\dots,\zeta_{B}$ independently injected into $\sqrt{V_{0}},\dots,\sqrt{V_{B}}$. The noise is scaled by $2\Delta/\xi$ where the global sensitivity satisfies $\Delta\lesssim 1/N$ for the MMD and HSIC square-rooted V-statistics (Kim and Schrab, 2023, Lemmas 5 and 6). The quantile $\widetilde{q}_{1-\alpha}$ of the privatised statistics is upper bounded by the sum of, the quantile of the statistics, and of the quantile of the Laplacian privatisation noise, we get
$\widetilde{q}_{1-\alpha}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C_{1}\sqrt{\frac{1}{N}\log\biggl{(}\frac{1}{\alpha}\biggr{)}}+C_{2}\frac{1}{N\xi}\log\biggl{(}\frac{1}{\alpha}\biggr{)},$ (66)
which holds with probability at least $31-\beta/$ 3. The above is derived by using a closed form on the cumulative distribution function (CDF) of the Laplace distribution $F^{-1}_{\zeta}(p)\coloneqq-\mathrm{sign}(p-0.5)\log(1-2|p-0.5|)$ for $p\in(0,1)$. Then, as before, we need
$\big{|}\sqrt{V_{0}}-{\mathrm{Kdisc}}\big{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C_{3}\sqrt{\frac{1}{N}\log\biggl{(}\frac{1}{\beta}\biggr{)}}$ (67)
to hold, this time with probability at least $31-\beta/$ 3, and also
$-\zeta_{0}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C_{4}\log\biggl{(}\frac{1}{\beta}\biggr{)}$ (68)
to hold with probability at least $31-\beta/$ 3 (again using the closed form of the Laplace CDF). Combining all these results we get
@@
${\mathrm{Kdisc}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \max\Biggl{\{}\sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}},\,\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N\xi}\Biggr{\}},$ (70)
@@ A.3 Proof sketch of robust kernel separation
We detail the proof structure of the robust kernel separation results: Equation 19 proved in Schrab and Kim (2025, Theorem 1.i), and Equation 34 proved in Schrab and Kim (2025, Theorem 2.i). We focus on the V-statistic case $\sqrt{V_{0}}$ with $V_{0}\coloneqq{V_{k}}$ for some kernel $k$, with bootstrapped statistics $\sqrt{V_{1}},\dots,\sqrt{V_{B}}$.
The proof structure follows closely the one outlined in Section A.1 but, for the robust tests, the quantile is shifted by a factor of $2r\Delta$ where $r$ is the robustness parameter and $\Delta$ is the global sensitivity, which for MMD and HSIC square-rooted V-statistics, scales as $1/N$ as shown in Kim and Schrab (2023, Lemmas 5 and 6). Adapting the reasoning of Section A.1, we then get
@@
${\mathrm{Kdisc}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \max\biggl{\{}\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta)\}}{N}},\ \frac{r}{N}\biggr{\}}$ (72)
@@ A.4 Proof sketch of pooled kernel separation
We detail the proof structure of the pooled kernel separation results: Equation 15 (fuse variant also proved in Biggs et al., 2023, Theorems 2 and 3) and Equations 30, 46, 29, 45 and 14. For simplicity, consider the tests to have square-rooted V-statistic $(T_{0})_{k}\coloneqq\sqrt{V_{k}}$ with kernel parameter $k\in\mathcal{K}$. We show that the type II error is guaranteed to be controlled by $\beta$ for all alternatives satisfying
$\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\leavevmode\nobreak\ {\mathrm{Kdisc}}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big{\}}}{N}}$ (73)
@@
$\max_{k\in{\mathcal{K}}}\,{\mathrm{Kdisc}}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big{\}}}{N}}$ (74)
for the pooled test over ${\mathcal{K}}$ with fuse or max pooling function. The uniform separation rate of Equation 73 holds trivially from the result for fixed kernel (Section A.1) and the fact the mean of discrepancies is equal to the discrepancy computed with a mean kernel (Schrab, 2025b, Equation 78), which is due to the linearity of the discrepancy in the kernel. For fuse and max kernel pooling, it is enough to prove that type II error control by $\beta$ is guaranteed when
$\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\leavevmode\nobreak\ {\mathrm{Kdisc}}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big{\}}}{N}},$ (75)
where ‘$\mathrm{pool}$’ is the corresponding pooling function (i.e., fuse or max). Indeed, when using the fuse pooling function, leveraging the relation between fuse and max (Schrab, 2025b, Equation 78), the ‘$\mathrm{pool}$’ function in Equation 75 can be replaced by ‘$\max$’ resulting in an additional additive term $\log(|{\mathcal{K}}|)/\nu$ term on the right hand side of Equation 75, which is absorbed in the rate provided that $\nu\geq\max(N,\log(|{\mathcal{K}}|))$ (as this implies $\log(|{\mathcal{K}}|)/\nu\leq\sqrt{\log(|{\mathcal{K}}|)/N}$).
Then, to prove Equation 75, following the reasoning of Section A.1, it suffices to show that
$\big{|}\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\leavevmode\nobreak\ (T_{0})_{k}-\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\leavevmode\nobreak\ {\mathrm{Kdisc}}_{k}\big{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C_{1}\sqrt{\frac{1}{N}\log\biggl{(}\frac{|{\mathcal{K}}|}{\beta}\biggr{)}}\qquad\textrm{ and }\qquad q_{1-\alpha}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C_{2}\sqrt{\frac{1}{N}\log\biggl{(}\frac{|{\mathcal{K}}|}{\alpha}\biggr{)}},$ (76)
each holding with probability at least $21-\beta/$ 2, where $q_{1-\alpha}$ is the quantile of the pooled bootstrapped statistics. Equivalently, as we are working with square-rooted V-statistics, we need to show that, for all $t>0$, we have
$\mathbb{P}\Big{(}\,\big{|}\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\leavevmode\nobreak\ (T_{0})_{k}-\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\leavevmode\nobreak\ {\mathrm{Kdisc}}_{k}\big{|}>\widetilde{t}\,\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ |{\mathcal{K}}|\exp(-C\,t^{2}N)$ (77)
@@
$\mathbb{P}\Big{(}\,\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\leavevmode\nobreak\ (T_{b})_{k}>\widetilde{t}\,\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ |{\mathcal{K}}|\exp(-\widetilde{C}\,t^{2}N),$ (78)
for the cases where ‘$\mathrm{pool}$’ is ‘fuse’, and is ‘max’, where $\widetilde{t}\coloneqq C^{\prime}N^{-1/2}+t$ as in Section A.1.
@@
As in Section A.1, each of the $|{\mathcal{K}}|$ single tests satisfies
$\big{|}(T_{0})_{k}-{\mathrm{Kdisc}}_{k}\big{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C_{k}\sqrt{\frac{1}{N}\log\biggl{(}\frac{|{\mathcal{K}}|}{\beta}\biggr{)}},$ (79)
@@
$\big{|}(T_{0})_{k}-{\mathrm{Kdisc}}_{k}\big{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C\sqrt{\frac{1}{N}\log\biggl{(}\frac{|{\mathcal{K}}|}{\beta}\biggr{)}},$ (80)
@@
$\big{|}\underset{k\in{\mathcal{K}}}{\mathrm{max}}\leavevmode\nobreak\ (T_{0})_{k}-\underset{k\in{\mathcal{K}}}{\mathrm{max}}\leavevmode\nobreak\ {\mathrm{Kdisc}}_{k}\big{|}\leavevmode\nobreak\ =\leavevmode\nobreak\ \big{|}(T_{0})_{k_{T}}-{\mathrm{Kdisc}}_{k_{K}}\big{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C\sqrt{\frac{1}{N}\log\biggl{(}\frac{|{\mathcal{K}}|}{\beta}\biggr{)}}.$ (81)
@@
This proves that Equation 81 holds.
@@
$\mathbb{P}\Big{(}\underset{k\in{\mathcal{K}}}{\mathrm{max}}\leavevmode\nobreak\ (T_{b})_{k}>\widetilde{t}\,\Big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \mathbb{P}\left(\bigcup_{k\in{\mathcal{K}}}\Big{\{}(T_{b})_{k}>\widetilde{t}\,\Big{\}}\right)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{k\in{\mathcal{K}}}\mathbb{P}\Big{(}(T_{b})_{k}>\widetilde{t}\,\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ |{\mathcal{K}}|\exp(-\widetilde{C}\,t^{2}N)$ (82)
for all $t>0$, where $\widetilde{C}=\min_{k\in{\mathcal{K}}}\widetilde{C}_{k}$, where $\mathbb{P}\big{(}(T_{b})_{k}>\widetilde{t}\,\big{)}\leq\exp(-\widetilde{C}_{k}\,t^{2}N)$ for all $t>0$ as in Section A.1, and where $\widetilde{t}\coloneqq C^{\prime}N^{-1/2}+t$.
@@
Again, as in Section A.1, starting from
$\big{|}(T_{0})_{k}-{\mathrm{Kdisc}}_{k}\big{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C\sqrt{\frac{1}{N}\log\biggl{(}\frac{|{\mathcal{K}}|}{\beta}\biggr{)}},$ (83)
@@
$\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\leavevmode\nobreak\ (T_{0})_{k}\leavevmode\nobreak$ $=\leavevmode\nobreak\ \frac{1}{\nu}\log\left(\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}\exp\big{(}\nu(T_{0})_{k}\big{)}\right)$ (84)
$\leq\leavevmode\nobreak\ \frac{1}{\nu}\log\left(\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}\exp\big{(}\nu({\mathrm{Kdisc}}_{k}+\delta)\big{)}\right)$
$=\leavevmode\nobreak\ \delta+\frac{1}{\nu}\log\left(\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}\exp\big{(}\nu{\mathrm{Kdisc}}_{k}\big{)}\right)$
$=\leavevmode\nobreak\ \delta+\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\leavevmode\nobreak\ {\mathrm{Kdisc}}_{k}$
@@
$\big{|}\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\leavevmode\nobreak\ (T_{0})_{k}-\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\leavevmode\nobreak\ {\mathrm{Kdisc}}_{k}\big{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C\sqrt{\frac{1}{N}\log\biggl{(}\frac{|{\mathcal{K}}|}{\beta}\biggr{)}}.$ (85)
@@
$\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\leavevmode\nobreak\ (T_{b})_{k}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \underset{k\in{\mathcal{K}}}{\mathrm{max}}\leavevmode\nobreak\ (T_{b})_{k}.$ (86)
@@
$\mathbb{P}\Big{(}\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\leavevmode\nobreak\ (T_{b})_{k}>\widetilde{t}\,\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathbb{P}\Big{(}\underset{k\in{\mathcal{K}}}{\mathrm{max}}\leavevmode\nobreak\ (T_{b})_{k}>\widetilde{t}\,\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ |{\mathcal{K}}|\exp(-\widetilde{C}\,t^{2}N)$ (87)
@@
A similar analysis can be done for the mean pooling function. While this is not necessary since we can get the uniform separation rate without the additional $\log(|{\mathcal{K}}|)$ term as discussed above in Equation 73, we present it nonetheless as this will be important for the global sensitivity discussion below.
@@
$\big{|}(T_{0})_{k}-{\mathrm{Kdisc}}_{k}\big{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C\sqrt{\frac{1}{N}\log\biggl{(}\frac{|{\mathcal{K}}|}{\beta}\biggr{)}},$ (88)
@@
$\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}(T_{0})_{k}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}\Big{(}{\mathrm{Kdisc}}_{k}+\delta\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \delta+\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}{\mathrm{Kdisc}}_{k}.$ (89)
@@
$\big{|}\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\leavevmode\nobreak\ (T_{0})_{k}-\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\leavevmode\nobreak\ {\mathrm{Kdisc}}_{k}\big{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C\sqrt{\frac{1}{N}\log\biggl{(}\frac{|{\mathcal{K}}|}{\beta}\biggr{)}}.$ (90)
The quantile bound follows as the one for the maximum case in Equation 82, that is
$\mathbb{P}\Big{(}\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\leavevmode\nobreak\ (T_{b})_{k}>\widetilde{t}\,\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathbb{P}\left(\bigcup_{k\in{\mathcal{K}}}\Big{\{}(T_{b})_{k}>\widetilde{t}\,\Big{\}}\right)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ |{\mathcal{K}}|\exp(-\widetilde{C}\,t^{2}N)$ (91)
@@
$\bigl{|}S_{k}(\mathbb{X}_{N}^{\pi})-S_{k}(\widetilde{\mathbb{X}}_{N}^{\pi})\bigr{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \Delta$ (92)

### v2

@@ A.2 Proof sketch of differentially private kernel separation
We detail the proof structure of the differentially private kernel separation results: Equation˜18 proved in Kim and Schrab (2023, Theorem 7), and Equation˜33 proved in Kim and Schrab (2023, Theorem 12). We focus on the V-statistic case $\sqrt{V_{0}}$ with $V_{0}\coloneqq{V_{k}}$ for some kernel $k$, with bootstrapped statistics $\sqrt{V_{1}},\dots,\sqrt{V_{B}}$.
The proof structure is exactly the same as outlined in Section˜A.1 but taking into account the Laplace privatisation noise $\zeta_{0},\dots,\zeta_{B}$ independently injected into $\sqrt{V_{0}},\dots,\sqrt{V_{B}}$. The noise is scaled by $2\Delta/\xi$ where the global sensitivity satisfies $\Delta\lesssim 1/N$ for the MMD and HSIC square-rooted V-statistics (Kim and Schrab, 2023, Lemmas 5 and 6). The quantile $\widetilde{q}_{1-\alpha}$ of the privatised statistics is upper bounded by the sum of, the quantile of the statistics, and of the quantile of the Laplacian privatisation noise, we get
$\widetilde{q}_{1-\alpha}\penalty 10000\ \leq\penalty 10000\ C_{1}\sqrt{\frac{1}{N}\log\biggl(\frac{1}{\alpha}\biggr)}+C_{2}\frac{1}{N\xi}\log\biggl(\frac{1}{\alpha}\biggr),$ (66)
which holds with probability at least $1-\beta/3$. The above is derived by using a closed form on the cumulative distribution function (CDF) of the Laplace distribution $F^{-1}_{\zeta}(p)\coloneqq-\mathrm{sign}(p-0.5)\log(1-2|p-0.5|)$ for $p\in(0,1)$. Then, as before, we need
$\big|\sqrt{V_{0}}-{\mathrm{Kdisc}}\big|\penalty 10000\ \leq\penalty 10000\ C_{3}\sqrt{\frac{1}{N}\log\biggl(\frac{1}{\beta}\biggr)}$ (67)
to hold, this time with probability at least $1-\beta/3$, and also
$-\zeta_{0}\penalty 10000\ \leq\penalty 10000\ C_{4}\log\biggl(\frac{1}{\beta}\biggr)$ (68)
to hold with probability at least $1-\beta/3$ (again using the closed form of the Laplace CDF). Combining all these results we get
@@
${\mathrm{Kdisc}}\penalty 10000\ \gtrsim\penalty 10000\ \max\Biggl\{\sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}},\,\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N\xi}\Biggr\},$ (70)
@@ A.3 Proof sketch of robust kernel separation
We detail the proof structure of the robust kernel separation results: Equation˜19 proved in Schrab and Kim (2025, Theorem 1.i), and Equation˜34 proved in Schrab and Kim (2025, Theorem 2.i). We focus on the V-statistic case $\sqrt{V_{0}}$ with $V_{0}\coloneqq{V_{k}}$ for some kernel $k$, with bootstrapped statistics $\sqrt{V_{1}},\dots,\sqrt{V_{B}}$.
The proof structure follows closely the one outlined in Section˜A.1 but, for the robust tests, the quantile is shifted by a factor of $2r\Delta$ where $r$ is the robustness parameter and $\Delta$ is the global sensitivity, which for MMD and HSIC square-rooted V-statistics, scales as $1/N$ as shown in Kim and Schrab (2023, Lemmas 5 and 6). Adapting the reasoning of Section˜A.1, we then get
@@
${\mathrm{Kdisc}}\penalty 10000\ \gtrsim\penalty 10000\ \max\biggl\{\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta)\}}{N}},\ \frac{r}{N}\biggr\}$ (72)
@@ A.4 Proof sketch of pooled kernel separation
We detail the proof structure of the pooled kernel separation results: Equation˜15 (fuse variant also proved in Biggs et al., 2023, Theorems 2 and 3) and Equations˜30, 46, 29, 45 and 14. For simplicity, consider the tests to have square-rooted V-statistic $(T_{0})_{k}\coloneqq\sqrt{V_{k}}$ with kernel parameter $k\in\mathcal{K}$. We show that the type II error is guaranteed to be controlled by $\beta$ for all alternatives satisfying
$\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\penalty 10000\ {\mathrm{Kdisc}}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big\}}{N}}$ (73)
@@
$\max_{k\in{\mathcal{K}}}\,{\mathrm{Kdisc}}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big\}}{N}}$ (74)
for the pooled test over $\mathcal{K}$} with fuse or max pooling function. The uniform separation rate of Equation˜73 holds trivially from the result for fixed kernel (Section˜A.1) and the fact the mean of discrepancies is equal to the discrepancy computed with a mean kernel (Schrab, 2025b, Equation 78), which is due to the linearity of the discrepancy in the kernel. For fuse and max kernel pooling, it is enough to prove that type II error control by $\beta$ is guaranteed when
$\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\penalty 10000\ {\mathrm{Kdisc}}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big\}}{N}},$ (75)
where $\mathrm{pool}$’ is the corresponding pooling function (i.e., fuse or max). Indeed, when using the fuse pooling function, leveraging the relation between fuse and max (Schrab, 2025b, Equation 78), the $\mathrm{pool}$’ function in Equation˜75 can be replaced by $\max$’ resulting in an additional additive term $\log(|{\mathcal{K}}|)/\nu$ term on the right hand side of Equation˜75, which is absorbed in the rate provided that $\nu\geq\max(N,\log(|{\mathcal{K}}|))$ (as this implies $\log(|{\mathcal{K}}|)/\nu\leq\sqrt{\log(|{\mathcal{K}}|)/N}$).
Then, to prove Equation˜75, following the reasoning of Section˜A.1, it suffices to show that
$\big|\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\penalty 10000\ (T_{0})_{k}-\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\penalty 10000\ {\mathrm{Kdisc}}_{k}\big|\penalty 10000\ \leq\penalty 10000\ C_{1}\sqrt{\frac{1}{N}\log\biggl(\frac{|{\mathcal{K}}|}{\beta}\biggr)}\qquad\textrm{ and }\qquad q_{1-\alpha}\penalty 10000\ \leq\penalty 10000\ C_{2}\sqrt{\frac{1}{N}\log\biggl(\frac{|{\mathcal{K}}|}{\alpha}\biggr)},$ (76)
each holding with probability at least $1-\beta/2$, where $q_{1-\alpha}$ is the quantile of the pooled bootstrapped statistics. Equivalently, as we are working with square-rooted V-statistics, we need to show that, for all $t>0$, we have
$\mathbb{P}\Big(\,\big|\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\penalty 10000\ (T_{0})_{k}-\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\penalty 10000\ {\mathrm{Kdisc}}_{k}\big|>\widetilde{t}\,\Big)\penalty 10000\ \leq\penalty 10000\ |{\mathcal{K}}|\exp(-C\,t^{2}N)$ (77)
@@
$\mathbb{P}\Big(\,\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\penalty 10000\ (T_{b})_{k}>\widetilde{t}\,\Big)\penalty 10000\ \leq\penalty 10000\ |{\mathcal{K}}|\exp(-\widetilde{C}\,t^{2}N),$ (78)
for the cases where $\mathrm{pool}$’ is ‘fuse’, and is ‘max’, where $\widetilde{t}\coloneqq C^{\prime}N^{-1/2}+t$ as in Section˜A.1.
@@
As in Section˜A.1, each of the $|{\mathcal{K}}|$ single tests satisfies
$\big|(T_{0})_{k}-{\mathrm{Kdisc}}_{k}\big|\penalty 10000\ \leq\penalty 10000\ C_{k}\sqrt{\frac{1}{N}\log\biggl(\frac{|{\mathcal{K}}|}{\beta}\biggr)},$ (79)
@@
$\big|(T_{0})_{k}-{\mathrm{Kdisc}}_{k}\big|\penalty 10000\ \leq\penalty 10000\ C\sqrt{\frac{1}{N}\log\biggl(\frac{|{\mathcal{K}}|}{\beta}\biggr)},$ (80)
@@
$\big|\underset{k\in{\mathcal{K}}}{\mathrm{max}}\penalty 10000\ (T_{0})_{k}-\underset{k\in{\mathcal{K}}}{\mathrm{max}}\penalty 10000\ {\mathrm{Kdisc}}_{k}\big|\penalty 10000\ =\penalty 10000\ \big|(T_{0})_{k_{T}}-{\mathrm{Kdisc}}_{k_{K}}\big|\penalty 10000\ \leq\penalty 10000\ C\sqrt{\frac{1}{N}\log\biggl(\frac{|{\mathcal{K}}|}{\beta}\biggr)}.$ (81)
@@
This proves that Equation˜81 holds.
@@
$\mathbb{P}\Big(\underset{k\in{\mathcal{K}}}{\mathrm{max}}\penalty 10000\ (T_{b})_{k}>\widetilde{t}\,\Big)\penalty 10000\ =\penalty 10000\ \mathbb{P}\left(\bigcup_{k\in{\mathcal{K}}}\Big\{(T_{b})_{k}>\widetilde{t}\,\Big\}\right)\penalty 10000\ \leq\penalty 10000\ \sum_{k\in{\mathcal{K}}}\mathbb{P}\Big((T_{b})_{k}>\widetilde{t}\,\Big)\penalty 10000\ \leq\penalty 10000\ |{\mathcal{K}}|\exp(-\widetilde{C}\,t^{2}N)$ (82)
for all $t>0$, where $\widetilde{C}=\min_{k\in{\mathcal{K}}}\widetilde{C}_{k}$, where $\mathbb{P}\big((T_{b})_{k}>\widetilde{t}\,\big)\leq\exp(-\widetilde{C}_{k}\,t^{2}N)$ for all $t>0$ as in Section˜A.1, and where $\widetilde{t}\coloneqq C^{\prime}N^{-1/2}+t$.
@@
Again, as in Section˜A.1, starting from
$\big|(T_{0})_{k}-{\mathrm{Kdisc}}_{k}\big|\penalty 10000\ \leq\penalty 10000\ C\sqrt{\frac{1}{N}\log\biggl(\frac{|{\mathcal{K}}|}{\beta}\biggr)},$ (83)
@@
[…]
$\leq\penalty 10000\ \frac{1}{\nu}\log\left(\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}\exp\big(\nu({\mathrm{Kdisc}}_{k}+\delta)\big)\right)$
$=\penalty 10000\ \delta+\frac{1}{\nu}\log\left(\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}\exp\big(\nu{\mathrm{Kdisc}}_{k}\big)\right)$
$=\penalty 10000\ \delta+\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\penalty 10000\ {\mathrm{Kdisc}}_{k}$
@@
$\big|\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\penalty 10000\ (T_{0})_{k}-\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\penalty 10000\ {\mathrm{Kdisc}}_{k}\big|\penalty 10000\ \leq\penalty 10000\ C\sqrt{\frac{1}{N}\log\biggl(\frac{|{\mathcal{K}}|}{\beta}\biggr)}.$ (85)
@@
$\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\penalty 10000\ (T_{b})_{k}\penalty 10000\ \leq\penalty 10000\ \underset{k\in{\mathcal{K}}}{\mathrm{max}}\penalty 10000\ (T_{b})_{k}.$ (86)
@@
$\mathbb{P}\Big(\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\penalty 10000\ (T_{b})_{k}>\widetilde{t}\,\Big)\penalty 10000\ \leq\penalty 10000\ \mathbb{P}\Big(\underset{k\in{\mathcal{K}}}{\mathrm{max}}\penalty 10000\ (T_{b})_{k}>\widetilde{t}\,\Big)\penalty 10000\ \leq\penalty 10000\ |{\mathcal{K}}|\exp(-\widetilde{C}\,t^{2}N)$ (87)
@@
A similar analysis can be done for the mean pooling function. While this is not necessary since we can get the uniform separation rate without the additional $\log(|{\mathcal{K}}|)$ term as discussed above in Equation˜73, we present it nonetheless as this will be important for the global sensitivity discussion below.
@@
$\big|(T_{0})_{k}-{\mathrm{Kdisc}}_{k}\big|\penalty 10000\ \leq\penalty 10000\ C\sqrt{\frac{1}{N}\log\biggl(\frac{|{\mathcal{K}}|}{\beta}\biggr)},$ (88)
@@
$\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}(T_{0})_{k}\penalty 10000\ \leq\penalty 10000\ \frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}\Big({\mathrm{Kdisc}}_{k}+\delta\Big)\penalty 10000\ \leq\penalty 10000\ \delta+\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}{\mathrm{Kdisc}}_{k}.$ (89)
@@
$\big|\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\penalty 10000\ (T_{0})_{k}-\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\penalty 10000\ {\mathrm{Kdisc}}_{k}\big|\penalty 10000\ \leq\penalty 10000\ C\sqrt{\frac{1}{N}\log\biggl(\frac{|{\mathcal{K}}|}{\beta}\biggr)}.$ (90)
The quantile bound follows as the one for the maximum case in Equation˜82, that is
$\mathbb{P}\Big(\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\penalty 10000\ (T_{b})_{k}>\widetilde{t}\,\Big)\penalty 10000\ \leq\penalty 10000\ \mathbb{P}\left(\bigcup_{k\in{\mathcal{K}}}\Big\{(T_{b})_{k}>\widetilde{t}\,\Big\}\right)\penalty 10000\ \leq\penalty 10000\ |{\mathcal{K}}|\exp(-\widetilde{C}\,t^{2}N)$ (91)
@@
$\bigl|S_{k}(\mathbb{X}_{N}^{\pi})-S_{k}(\widetilde{\mathbb{X}}_{N}^{\pi})\bigr|\penalty 10000\ \leq\penalty 10000\ \Delta$ (92)